A consumer organization inspecting new cars found that many had appearance defects​ (dents, scratches, paint​ chips, etc.). While none had more than three of these​ defects, 9​% had​ three, 15​% had​ two, and 28​% had one defect. Find the expected number of appearance defects in a new car and the standard deviation.
Answer #1

from given data :

x f(x) xP(x) x2P(x)
0 0.4800 0.0000 0.0000
1 0.2800 0.2800 0.2800
2 0.1500 0.3000 0.6000
3 0.0900 0.2700 0.8100
total 0.8500 1.6900
E(x) =μ= ΣxP(x) = 0.8500
E(x2) = Σx2P(x) = 1.6900
Var(x)=σ2 = E(x2)-(E(x))2= 0.9675
std deviation=         σ= √σ2 = 0.9836

from above

expected number of appearance defects =0.85

and  standard deviation =0.9836